How does the answer equal -8 in the equation 20-(8+x)=-(1-x)+29?

Answer by longjonsilver(2297)

20-(8+x)=-(1-x)+29?
20-8-x = -1+x+29
12-x = x+28
12 = 2x + 28
or
2x + 28 = 12
2x = -16
x = -8

jon.